Last month you assumed the position of manager for a large car dealership. The distinguishing feature of this dealership is its “no hassle” pricing strategy; prices (usually well below sticker) are posted on the windows, and your sales staff has a reputation for not negotiating with customers. Last year, the company spent $2 million on advertisements to inform customers about its “no hassle” policy and had revenues of $40 million. A recent study from an ad agency indicates that for each 3 percent increase in TV advertising expenditures, a car dealer can expect to sell 12 percent more cars – but that it would take a 4 percent decrease in price to generate the same 12 percent increase in units sole. Assuming the information is correct, should you increase or decrease your firm’s level of advertising? Explain. Show all calculations and provide the reasoning clearly and concise as possible.
Answer #1

Advertising elasticity of demand = % change in D / % change in A

Advertising elasticity of demand = 12/3

Advertising elasticity of demand = 4

Similarly,

Price elasticity of demand = 12/4

Price elasticity of demand = 3

Now,

One can find the optimal profit maximizing A-sales ratio as below:

A/R = Advertising elasticity / Price elasticity

A/R = 4/3

A/R = 1.33

This means the firm must spend 1.33% of total sales of 40 million = $5.32 million on advertising

However, currently it is spending only $2 million on advertising

Thus, it is optimal for the firm to increase its level of advertising in order to reach its optimal level.